# Unveiling the impact of cross-order hyperdegree correlations in contagion processes on hypergraphs

### v1

@@ V Results
Gillespie simulations were performed on these higher-order hypernetworks to compare with our model predictions. To facilitate the analysis, we define the effective infection rates $\lambda_{1}=\langle k_{1}\rangle\beta_{1}/\mu$ and $\lambda_{2}=\langle k_{2}\rangle\beta_{2}/\mu$. Figure 2 shows the comparison between the compact effective hyperdegree model and simulations for both the temporal evolution of the proportion of infected nodes and the phase diagram of prevalence (i.e., the final proportion of infected individuals) as a function of $\lambda_{1}$.
Figure 2: EHDM vs simulation comparison: We compare the model against stochastic simulations for three different classes of hypergraphs: a regular one where all nodes have $k_{1}=5$ and $k_{2}=3$; a hypergraph with Poisson hyperdegree distribution with average $\langle k_{1}\rangle=5$ and $\langle k_{2}\rangle=3$; lastly, a hypergraph exhibiting a truncated power-law hyperdegree distribution with characteristic exponent $\nu_{1}=2.5$ for pairwise interactions and $\nu_{2}=2.25$ for three-body interactions. We consider $N=1000$, and a recovery rate $\gamma=1$ for all cases. Panels (a–c) compare the temporal evolution of the proportion of infected individuals between the compact effective hyperdegree model (red line) and Gillespie simulations (orange lines). For these cases $\lambda_{1}=1.5$ and $\lambda_{2}=3$. Panels (d–f) show the phase diagram of the final proportion of infected individuals $\rho$ as a function of $\lambda_{1}$, showing the bistable regime. In these cases $\lambda_{2}=2.5$.
@@ V.1 Effect of hyperdegree heterogeneity
Figure 3: Role of hyperdegree heterogeneity: (a) shows the final epidemic size from both stochastic simulations (diamonds) and the solution of the compact effective hyperdegree model (continuous line) as a function of the effective infection rate $\lambda_{1}$ for five different hypernetworks, all exhibiting negative binomial hyperdegree distributions for both pairs and triples with $N=1000$, $\langle k_{1}\rangle\approx 5$, $\langle k_{1}\rangle\approx 3$. The structures differ in the variance of their hyperdegree distributions, taking values $10$, $30$, $60$, $100$, and $300$. In panels (b–e), we display heat-maps of the difference in prevalence $\Delta\rho$ between two epidemic processes on hypernetworks with the same properties and the same infection and recovery rates, but different initial numbers of infected individuals. Each panel corresponds to a higher-order network with increasing variance $10$, $60$ and $80$.
@@ V.2 Effect of cross-order hyperdegree correlations
Figure 4: Effect of cross-order hyperdegree correlation: (a) displays the final epidemic size from both simulations (diamonds) and the solution of the compact effective hyperdegree model (solid lines) as a function of the infection rate $\lambda_{1}$ for heterogeneous hypernetworks with varying cross-order hyperdegree correlation, from $\sigma=-1$ to $\sigma=1$ in different colors, for all curves $\lambda_{2}=3$. In panel (b), we show the value of the critical infection rate $\lambda_{1}^{*}$ found through the EDHO model as a function of cross-order hyperdegree correlation for higher-order networks with different levels of heterogeneity. Panel (c) showcases the difference between temporal centroids of pairwise infections $I_{PW}$ and group infections $I_{HO}$ as a function of cross-order hyperdegree correlations for higher-order networks with different levels of heterogeneity. Panels (d–h) display the hyperdegree distributions of hypernetworks with a different cross-order hyperdegree correlation, specifically $\sigma=-1,-0.5,0,0.5,1$. Panels (i–m) show the temporal evolution of the number of infected individuals II and the pairwise and group contributions $I_{\rm PW}$ and $I_{\rm HO}$ for simulations on hypernetworks with hyperdegree distributions shown in panels (d–h). All hypergraphs have size $N=1000$ and exhibit negative binomial hyperdegree distributions for both pairwise and triple interactions, with average hyperdegrees $\langle k_{1}\rangle\approx 6$ and $\langle k_{2}\rangle\approx 6$, and a hyperdegree variance of approximately 30 for both interaction orders (except in panel (c), where the variance differs for each curve as indicated in the legend).
@@
$\begin{split}&\dot{I}_{\rm PW}=\sum_{k_{1},k_{2}}\bigg(\beta_{1}k_{1}\langle I\rangle S_{k_{1}}^{k_{2}}\bigg)-\mu I_{\rm PW}\\ &\dot{I}_{\rm HO}=\sum_{k_{1},k_{2}}\bigg(\beta_{2}k_{2}\langle Z\rangle S_{k_{1}}^{k_{2}}\bigg)-\mu I_{\rm HO}\end{split}$ (7)
@@ V.3 Hierarchical Spread in higher-order networks
Figure 5: Hierarchical Spread in Higher-Order Networks: In panel (a), we show the temporal evolution of the average 1-hyperedge degree of newly infected nodes, and in panel (b) the corresponding average 2-hyperedge degree. Panels (c) and (d) display the time evolution of the participation ratio for pairwise and group interactions, respectively, as defined in (10). All hypergraphs have size $N=1000$ and exhibit negative binomial hyperdegree distributions for both pairwise and three-body interactions. In all four panels, each curve corresponds to a distinct value of the cross-order hyperdegree correlation, $\sigma\in{-1,-0.5,0,0.5,1}$. The hyperdegree distributions used in the simulations have mean values $\langle k_{1}\rangle\approx 6$ and $\langle k_{2}\rangle\approx 6$, with a variance of 30 for both interaction orders. Disease parameters are fixed at $\lambda_{1}=0.9$ and $\lambda_{2}=3$.
@@ V.4 Control of influential nodes
Figure 6: Control of influential nodes: Panels a) to e) show the results of the prevalence reduction $\Delta\varrho$ as a function of the control parameter $\alpha$. Each panel has a different value of $\lambda_{2}$. In all four panels, each curve corresponds to a distinct value of the cross-order hyperdegree correlation, $\sigma\in{-1,-0.5,0,0.5,1}$. The hyperdegree distributions used in the simulations have mean values $\langle k_{1}\rangle\approx 6$ and $\langle k_{2}\rangle\approx 6$, with a variance of 30 for both interaction orders. The disease parameters were set to $\lambda_{1}=0.9$, while $\lambda_{2}$ was varied over the values $[0,1,2,5,10]$.

### v2

@@ V Results
Gillespie simulations were performed on these higher-order hypernetworks to compare with our model predictions. To facilitate the analysis, we define the effective infection rates $\lambda_{1}=\langle k_{1}\rangle\beta_{1}/\gamma$ and $\lambda_{2}=\langle k_{2}\rangle\beta_{2}/\gamma$. Figure 2 shows the comparison between the compact effective hyperdegree model and simulations for both the temporal evolution of the proportion of infected nodes and the phase diagram of prevalence (i.e., the final proportion of infected individuals) as a function of $\lambda_{1}$.
Figure 2: EHDM vs simulation comparison: We compare the effective hyperdegree model against stochastic simulations across three distinct hypergraph classes: a regular hypergraph where all nodes have $k_{1}=5$ and $k_{2}=3$; a hypergraph with Poisson-distributed hyperdegrees with means $\langle k_{1}\rangle=5$ and $\langle k_{2}\rangle=3$; and a hypergraph exhibiting truncated power-law hyperdegree distributions with characteristic exponents $\nu_{1}=2.5$ for pairwise interactions and $\nu_{2}=2.25$ for three-body interactions. All simulations use $N=1000$ nodes and a recovery rate $\gamma=1$. Panels (a–c) compare the temporal evolution of the infected population fraction between the effective hyperdegree model (red line) and Gillespie simulations (orange lines), with infection rates $\lambda_{1}=1.5$ and $\lambda_{2}=3$. Panels (d–f) show phase diagrams of the final infected fraction $\rho$ as a function of $\lambda_{1}$, comparing stochastic simulations (diamonds and circles) with the EHDM (solid lines). In these cases we set $\lambda_{2}=2.5$, and different colors indicate different initial numbers of infected nodes, and indicated in the legend.
@@ V.1 Effect of hyperdegree heterogeneity
Figure 3: Role of hyperdegree heterogeneity: Panel (a) shows the final epidemic size as a function of the effective infection rate $\lambda_{1}$ for five different hypernetworks, comparing stochastic simulations (diamonds) with the effective hyperdegree model (solid lines). All networks have $N=1000$ nodes with negative binomial hyperdegree distributions for both pairs and triples, average hyperdegrees $\langle k_{1}\rangle\approx 5$ and $\langle k_{2}\rangle\approx 3$, and varying hyperdegree distribution variance (10, 30, 60, 100, and 300). Panels (b–d) display heat-maps of the prevalence difference $\Delta\rho$ between two epidemic processes with identical infection and recovery rates but different initial numbers of infected individuals, for networks with variance 10, 60, and 80. The recovery rate is set to $\gamma=1$ for all cases.
@@ V.2 Effect of cross-order hyperdegree correlations
Figure 4: Effect of cross-order hyperdegree correlation: Panel (a) shows the final epidemic size as a function of infection rate $\lambda_{1}$ for heterogeneous hypernetworks with varying cross-order hyperdegree correlations ($\sigma=-1$ to $1$, shown in different colors), where stochastic simulations (diamonds) are compared with the effective hyperdegree model (solid lines) for $\lambda_{2}=3$. Panel (b) presents the bistability index $\mathcal{B}$ as a function of $\lambda_{2}$ across the same range of correlation values for $\lambda_{1}=0.9$. Panel (c) illustrates the difference between temporal centroids of pairwise infections ($I_{\text{PW}}$) and group infections ($I_{\text{HO}}$) as a function of $\sigma$ for five networks with different heterogeneity levels, quantified by the hyperdegree distribution variance indicated in the legend. Panels (d–h) display the hyperdegree distributions for networks with $\sigma=-1,-0.5,0,0.5$, and $1$, respectively. Panels (i–m) show the temporal evolution of the total infected fraction II, along with the contributions from pairwise ($I_{\text{PW}}$) and group ($I_{\text{HO}}$) infections, for hypernetworks with cross-order correlations corresponding to the hyperdegree distributions shown in panels (d–h). All hypernetworks contain $N=1000$ nodes with negative binomial hyperdegree distributions for both pairwise and triple interactions, average hyperdegrees $\langle k_{1}\rangle\approx\langle k_{2}\rangle\approx 6$, and variance $\approx 30$ for both interaction orders, except in panel (c) where variance varies as indicated in the legend. Disease parameters for panels (c) and (i–m) are $\lambda_{1}=0.9$, $\lambda_{2}=3$, and $\gamma=1$ for all cases.
@@
$\begin{split}&\dot{I}_{\rm PW}=\sum_{k_{1},k_{2}}\bigg(\beta_{1}k_{1}\langle I\rangle S_{k_{1}}^{k_{2}}\bigg)-\gamma I_{\rm PW}\\ &\dot{I}_{\rm HO}=\sum_{k_{1},k_{2}}\bigg(\beta_{2}k_{2}\langle Z\rangle S_{k_{1}}^{k_{2}}\bigg)-\gamma I_{\rm HO}\end{split}$ (7)
@@ V.3 Hierarchical Spread in higher-order networks
Figure 5: Hierarchical Spread in Higher-Order Networks:Panels (a) and (b) show the temporal evolution of the average 1-hyperedge degree and 2-hyperedge degree, respectively, of newly infected nodes. Panels (c) and (d) display the corresponding time evolution of the participation ratio for pairwise and group interactions, as defined in (10). In all panels, each curve corresponds to a distinct value of the cross-order hyperdegree correlation, $\sigma=-1,-0.5,0,0.5$, and $1$. All hypergraphs have $N=1000$ nodes with negative binomial hyperdegree distributions for both pairwise and three-body interactions, mean hyperdegrees $\langle k_{1}\rangle\approx\langle k_{2}\rangle\approx 6$, and variance 30 for both interaction orders. Disease parameters are $\lambda_{1}=0.9$, $\lambda_{2}=3$ and $\gamma=1$.
@@ V.4 Control of influential nodes
Figure 6: Control of influential nodes: Panels (a–e) show the prevalence reduction $\Delta\varrho$ as a function of the control parameter $\alpha$ for different values of $\lambda_{2}=0,1,2,5$ and $10$. In all panels, each curve corresponds to a distinct value of the cross-order hyperdegree correlation, $\sigma=-1,-0.5,0,0.5$ and $1$. All hypergraphs have negative binomial hyperdegree distributions for both pairwise and three-body interactions with mean hyperdegrees $\langle k_{1}\rangle\approx\langle k_{2}\rangle\approx 6$ and variance 30 for both interaction orders. The infection rate for pairwise interactions is $\lambda_{1}=0.9$.